### v2

@@ Appendix A Proofs
$\mathbb{E}\left[\left.\left|D_{32}\right|^{2}\right|{\bf X}\right]$ $=n^{-1}\left(\widehat{H}_{2}-H_{2}\right)r^{{}^{\prime}}\mathbb{E}\left[\left.\xi\xi^{{}^{\prime}}\right|{\bf X}\right]r\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}$
$\leq C\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}$
@@
$=O_{p}\left(n^{-1}\zeta\left(M_{3}\right)^{2}M_{3}\right)=o_{p}\left(1\right),$
@@
$\left|D_{33}\right|^{2}$ $\leq n\left(\widehat{H}_{2}\widehat{R}\widehat{H}_{2}^{{}^{\prime}}\right)n^{-1}\sum_{i}\left(b\left(x_{i}\right)-r_{i}^{{}^{\prime}}\rho^{M_{3}}\right)^{2}$
$=O_{p}\left(nM_{3}^{-2d_{3}/j_{3}}\right)=o_{p}\left(1\right),$
@@
$\left|D_{34}\right|^{2}=nF^{2}\left(b\left(x\right)-r\left(x\right)^{{}^{\prime}}\rho^{M_{3}}\right)^{2}=O_{p}\left(nM_{3}^{-2d_{3}/j_{3}}\right)=o_{p}\left(1\right).$
@@
$n^{1/2}F\left(\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)=n^{-1/2}\sum_{i}H_{2}r_{i}\xi_{i}+o_{p}\left(1\right).$ (A.40)
@@
$n^{1/2}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-a\left(b_{1},V\right)\right)$ $=n^{-1/2}\sum_{i}\left(\psi_{1i}+\psi_{2i}+\psi_{3i}\right)+o_{p}\left(1\right),$ (A.41)
@@
Furthermore, notice that
@@
$\leq$ $\ Cn^{-1}\left(\zeta\left(M_{2}\right)^{2}M_{2}+\zeta\left(M_{2}\right)^{4}\zeta\left(M_{1}\right)^{4}M_{1}+\zeta\left(M_{3}\right)^{4}\zeta\left(M_{1}\right)^{4}M_{1}+\zeta\left(M_{3}\right)^{2}M_{3}\right)\rightarrow 0,$
where the last inequality holds by Lemma B5 of IN02. Then, by the Lindeberg–Feller CLT,
[…]
where $\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)=\widehat{b}\left(x\right)$ and $a\left(b_{1},V\right)=b\left(x\right)$.
@@
$\widehat{\Omega}_{2}$ $\coloneqq\widehat{\Omega}_{21}+\widehat{\Omega}_{22},\ \widehat{\Omega}_{21}\coloneqq\widehat{A}_{1}\widehat{P}^{-1}\left(n^{-1}\sum_{i}\widehat{p}_{i}\widehat{p}_{i}^{{}^{\prime}}\widehat{u}_{i}^{2}\right)\widehat{P}^{-1}\widehat{A}_{1}^{{}^{\prime}},$ (A.45)
$\widehat{\Omega}_{22}$ $\coloneqq n^{-1}\sum_{i}\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-r^{M_{3}}\left(x\right)^{{}^{\prime}}\left(\widehat{\overline{\mu}}_{i}^{II}+r_{i}\widehat{\xi}_{i}\right)\right)\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-r^{M_{3}}\left(x\right)^{{}^{\prime}}\left(\widehat{\overline{\mu}}_{i}^{II}+r_{i}\widehat{\xi}_{i}\right)\right)^{{}^{\prime}},$
$\widehat{A}_{1}$ $\coloneqq r^{M_{3}}\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}\left(n^{-1}\sum_{i}r_{i}\widehat{\overline{p}}_{i}^{{}^{\prime}}\right),\ \text{$\widehat{\overline{p}}_{i}$}\coloneqq\overline{p}\left(\widehat{s}_{i}\right),$
$\widehat{\overline{\mu}}_{i}^{II}$ $\coloneqq n^{-1}\sum_{j}r_{j}\widehat{b}_{1,V}\left(\widehat{v}_{j},w_{j}\right)q_{j}^{{}^{\prime}}\widehat{Q}^{-1}q_{i}\widehat{v}_{ji},\text{and }\widehat{\xi}_{i}\coloneqq\widehat{b}_{1}\left(\widehat{v}_{i},w_{i}\right)-\widehat{b}\left(x_{i}\right).$
[…]
@@
[…]
@@
$\leq C\zeta\left(M_{3}\right)^{2}n^{-1}\sum_{i}\left[\left(\widehat{\overline{p}}_{i}^{{}^{\prime}}\left(\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right)\right)^{2}+\left(\widehat{\overline{p}}_{i}^{{}^{\prime}}\alpha^{M_{2}}-b_{1}\left(\widehat{v}_{i},w_{i}\right)\right)^{2}\right]$
@@
$\mathbb{E}\left[\left.n^{-1}\sum_{i}D_{44i}^{2}\right|{\bf X}\right]$ $=\left(\widehat{H}_{2}-H_{2}\right)n^{-1}\sum_{i}r_{i}r_{i}^{{}^{\prime}}\mathbb{E}\left(\left.\xi_{i}^{2}\right|X_{i}\right)\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}$
$\leq C\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}$
@@
Combining the results for $D_{41}$, $D_{42}$, $D_{43}$, and $D_{44}$, I obtain
@@
$+3n^{-1}\sum_{i}\left(\widehat{H}_{2}r_{i}\widehat{\xi}_{i}-H_{2}r_{i}\xi_{i}\right)^{2}=o_{p}\left(1\right).$
@@
[…]
For the last result of Theorem 3, first define
$\widehat{\Omega}_{3}$ $\coloneqq\widehat{\Omega}_{31}+\widehat{\Omega}_{32},\ \widehat{\Omega}_{31}\coloneqq\widehat{A}_{3}\widehat{P}^{-1}\left(n^{-1}\sum_{i}\widehat{p}_{i}\widehat{p}_{i}^{{}^{\prime}}\widehat{u}_{i}^{2}\right)\widehat{P}^{-1}\widehat{A}_{3},$ (A.58)
$\widehat{\Omega}_{32}$ $\coloneqq n^{-1}\sum_{i}\left(\widehat{A}_{3}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-\left(\widehat{\overline{\mu}}_{i}^{III}+\widehat{\xi}_{i}\right)\right)\left(\widehat{A}_{3}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-\left(\widehat{\overline{\mu}}_{i}^{III}+\widehat{\xi}_{i}\right)\right)^{{}^{\prime}},$
$\widehat{A}_{3}$ $\coloneqq n^{-1}\sum_{i}\widehat{\overline{p}}_{i}^{{}^{\prime}},\text{ and}\ \widehat{\overline{\mu}}_{i}^{III}\coloneqq n^{-1}\sum_{j}\widehat{b}_{1,V}\left(\widehat{v}_{j},w_{j}\right)q_{j}^{{}^{\prime}}\widehat{Q}^{-1}q_{i}\widehat{v}_{ji}.$
[…]
Appendix B Sufficient Conditions for Assumption 2
Proposition 1 (Sufficient Conditions for Assumption 2 based on Altonji and Matzkin (2005)).
@@ Appendix B Sufficient Conditions for Assumption 2
To begin with, we use the first condition $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ to establish the exchangeability condition
$f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}=f_{\left.A_{i}\right|\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}},$ (B.1)
where $\left(\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}\right)$ is any permutation of $\left({\bf X}_{i},{\bf Z}_{i}\right)$ in time. It is worth emphasizing that (B.1), which will be proved in the following, is different from Assumption 2.3 of Altonji and Matzkin (2005). [^27] To be specific, Altonji and Matzkin (2005) assume an exchangeability condition involving both $A$ and $\eta_{t}$, i.e., $f_{\left.A_{i},\eta_{it}\right|{\bf X}_{i}}=f_{\left.A_{i},\eta_{it}\right|{\bf\widetilde{X}}_{i}},$ which effectively rules out time-varying endogeneity through the random coefficients since it requires that the density of $\eta_{it}$ given $X_{it}=x_{it}$ is the same as that given $X_{it}=x_{is}$ for any $s\neq t$. Without loss of generality, we prove (B.1) for $T=2$ since any ordering of $\left\{1,...,T\right\}$ for finite $T$ can be achieved via a finite number of pairwise permutations. Then, we prove that one can construct $W_{i}$ such that Assumption 2 holds. For simplicity of notation, we assume $X_{it}$ and $Z_{it}$ are both scalars and suppress $i$ subscripts in all variables. Thus, in this proof all subscripts of the variables denote the time period.
@@
$f_{A,\eta_{1},\eta_{2}}\left(a,h_{1},h_{2}\right)=f_{A,\eta_{1},\eta_{2}}\left(a,h_{2},h_{1}\right).$ (B.2)
@@
$=$ $\ f_{\left.X_{1},X_{2},A\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1},a\right|z_{2},z_{1}\right),$ (B.3)
where the first equality holds by the change of variables for $\eta_{1}$ and $\eta_{2}$, the second equality uses (B.2) and $Z\perp\left(A,\eta\right)$, the third equality holds by $X_{1}=g\left(z_{2},a,g^{-1}\left(x_{2},z_{2},a\right)\right)=x_{2}$ and $X_{2}=g\left(z_{1},a,g^{-1}\left(x_{1},z_{1},a\right)\right)=x_{1}$, and the last equality uses the fact that the product of derivatives of inverse functions equals one.
Given (B.3), I integrate $A$ out in (B.3) to obtain
$f_{\left.X_{1},X_{2}\right|Z_{1},Z_{2}}\left(\left.x_{1},x_{2}\right|z_{1},z_{2}\right)$ $=f_{\left.X_{1},X_{2}\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1}\right|z_{2},z_{1}\right),$ (B.4)
@@
$=$ $\ f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{2},x_{1},z_{2},z_{1}\right),$ (B.5)
where the second equality uses (B.3) and (B.4).
The rest of the proof follows the same argument as in Section 2.2 of Altonji and Matzkin (2005). Specifically, I show that for any on-support $a$, the conditional density $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by a function of the form $f_{\left.A\right|W}\left(\left.a\right|w\right)$, where $W$ is a vector-valued function symmetric in the paired elements of $\left(X,Z\right)$. See footnote 10 for the detailed description of $W$. By condition (b) of Proposition 1, the supports of $X$ and $Z$ are compact. By condition (c) of Proposition 1, $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}$ is continuous in $\left(X_{1},X_{2},Z_{1},Z_{2}\right)$. Therefore, by the Stone-Weierstrass theorem, there exists a function $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}$ that is a polynomial in $\left(X_{1},X_{2},Z_{1},Z_{2}\right)$ over a compact set with the property that for any fixed $\delta$ that is arbitrarily close to 0,
$\max_{x_{1},x_{2}\in\mathcal{X},z_{1},z_{2}\in\mathcal{Z}}\left|f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)-f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)\right|\leq\delta.$ (B.6)
@@
denote the simple average of $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ over all $T!$ (here $T=2$) unique permutations of $\left(x_{t},z_{t}\right)$, and similarly for $\overline{f}_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$. By (B.5),

### v3

@@ A.1 Proofs
$\mathbb{E}\left[\left.\left|D_{32}\right|^{2}\right|{\bf X}\right]$ $=n^{-1}\left(\widehat{H}_{2}-H_{2}\right)r^{\prime}\mathbb{E}\left[\left.\xi\xi^{\prime}\right|{\bf X}\right]r\left(\widehat{H}_{2}-H_{2}\right)^{\prime}$
$\leq C\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{\prime}$
@@
$=O_{p}\left(n^{-1}\zeta\left(M_{3}\right)^{2}M_{3}\right)=o_{p}\left(1\right),$ (A.39)
@@
$\left|D_{33}\right|^{2}$ $\leq n\left(\widehat{H}_{2}\widehat{R}\widehat{H}_{2}^{\prime}\right)n^{-1}\sum_{i}\left(b\left(x_{i}\right)-r_{i}^{\prime}\rho^{M_{3}}\right)^{2}$
$=O_{p}\left(nM_{3}^{-2d_{3}/j_{3}}\right)=o_{p}\left(1\right),$ (A.40)
@@
$\left|D_{34}\right|^{2}=nF^{2}\left(b\left(x\right)-r\left(x\right)^{\prime}\rho^{M_{3}}\right)^{2}=O_{p}\left(nM_{3}^{-2d_{3}/j_{3}}\right)=o_{p}\left(1\right).$ (A.41)
@@
$\sqrt{n}F\left(\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)=n^{-1/2}\sum_{i}H_{2}r_{i}\xi_{i}+o_{p}\left(1\right).$ (A.42)
@@
$\sqrt{n}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-a\left(b_{1},V\right)\right)$ $=n^{-1/2}\sum_{i}\left(\psi_{1i}+\psi_{2i}+\psi_{3i}\right)+o_{p}\left(1\right),$ (A.43)
@@
Furthermore, observe that
@@
$\leq$ $\ Cn^{-1}\left(\zeta\left(M_{2}\right)^{2}M_{2}+\zeta\left(M_{2}\right)^{4}\zeta\left(M_{1}\right)^{4}M_{1}+\zeta\left(M_{3}\right)^{4}\zeta\left(M_{1}\right)^{4}M_{1}+\zeta\left(M_{3}\right)^{2}M_{3}\right)\rightarrow 0,$ (A.46)
where the last inequality holds by Lemma B5 of IN02.
Then, by the Lindeberg–Feller CLT,
$\sqrt{n}\Omega_{2}^{-1/2}\left(\widehat{b}\left(x\right)-b\left(x\right)\right)\overset{d}{\longrightarrow}N\left(0,1\right).$ (A.47)
@@
$\widehat{\Omega}_{2}$ $\coloneqq\widehat{\Omega}_{21}+\widehat{\Omega}_{22},\text{ where}$
$\widehat{\Omega}_{21}$ $\coloneqq\widehat{A}_{1}\widehat{P}^{-1}\left(n^{-1}\sum_{i}\widehat{p}_{i}\widehat{p}_{i}^{\prime}\widehat{u}_{i}^{2}\right)\widehat{P}^{-1}\widehat{A}_{1}^{\prime},$
$\widehat{\Omega}_{22}$ $\coloneqq n^{-1}\sum_{i}\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-r\left(x\right)^{\prime}\widehat{R}^{-1}\left(\widehat{\overline{\mu}}_{i}^{II}+r_{i}\widehat{\xi}_{i}\right)\right)\left(\widehat{A}_{1}\widehat{P}^{-1}\widehat{\overline{\mu}}_{i}^{I}-r\left(x\right)^{\prime}\widehat{R}^{-1}\left(\widehat{\overline{\mu}}_{i}^{II}+r_{i}\widehat{\xi}_{i}\right)\right)^{\prime},$
$\widehat{A}_{1}$ $\coloneqq r\left(x\right)^{\prime}\widehat{R}^{-1}\left(n^{-1}\sum_{i}r_{i}\overline{p}\left(\widehat{s}_{i}\right)^{\prime}\right),$
$\widehat{\overline{\mu}}_{i}^{II}$ $\coloneqq n^{-1}\sum_{j}r_{j}\widehat{b}_{1,V}\left(\widehat{v}_{j},w_{j}\right)q_{j}^{\prime}\widehat{Q}^{-1}q_{i}\widehat{v}_{ji},\text{ and }$
$\widehat{\xi}_{i}$ $\coloneqq\widehat{b}_{1}\left(\widehat{v}_{i},w_{i}\right)-\widehat{b}\left(x_{i}\right).$ (A.48)
I show $\widehat{\Omega}_{2}/\Omega_{2}-1\overset{p}{\longrightarrow}0$.
Recall that
@@
The proof of $\widehat{\Omega}_{21}/\Omega_{2}-\Omega_{21}/\Omega_{2}\overset{p}{\longrightarrow}0$ is almost identical to the proof of Lemma B10 of IN02, except that $\widehat{A}_{1}$ instead of $A_{1}$ appears in the definition of $\widehat{H}_{1}$. Nonetheless, I have shown $\left\|\widehat{H}_{1}-H_{1}\right\|=o_{p}\left(1\right)$. Thus, the proof of Lemma B10 of IN02 directly applies.
@@
$\leq C\zeta\left(M_{3}\right)^{2}n^{-1}\sum_{i}\left[\left(\widehat{\overline{p}}_{i}^{\prime}\left(\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right)\right)^{2}+\left(\widehat{\overline{p}}_{i}^{\prime}\alpha^{M_{2}}-b_{1}\left(\widehat{v}_{i},w_{i}\right)\right)^{2}\right]$
@@
$\mathbb{E}\left[\left.n^{-1}\sum_{i}D_{44i}^{2}\right|{\bf X}\right]$ $=\left(\widehat{H}_{2}-H_{2}\right)n^{-1}\sum_{i}r_{i}r_{i}^{\prime}\mathbb{E}\left(\left.\xi_{i}^{2}\right|X_{i}\right)\left(\widehat{H}_{2}-H_{2}\right)^{\prime}$
$\leq C\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{\prime}$
@@
Combining the results for $D_{41}$, $D_{42}$, $D_{43}$, and $D_{44}$, I have
@@
$+3n^{-1}\sum_{i}\left(\widehat{H}_{2}r_{i}\widehat{\xi}_{i}-H_{2}r_{i}\xi_{i}\right)^{2}=o_{p}\left(1\right).$ (A.58)
@@
$\widehat{\Omega}_{2}/\Omega_{2}-1\overset{p}{\longrightarrow}0.$ (A.62)
Next, for the APE estimator $\widehat{\overline{b}}$, define $b_{t}^{\text{APE}}=\mathbb{E}\beta_{it}$ and thus $\overline{b}=T^{-1}\sum_{t=1}^{T}b_{t}^{\text{APE}}$. I first estimate $b_{t}^{\text{APE}}$ by setting $r^{M_{3}}(x)=1$ in (4.7) for $\widehat{b}_{t}(x)$. Hence, the asymptotic normality proof for $\widehat{b}_{t}(x)$ directly applies to $\widehat{b}_{t}^{\text{APE}}$. In particular, by adapting to (A.43) with $r^{M_{3}}(x)=1$, the influence function for $\widehat{b}_{t}^{\text{APE}}$ in period $t$ is
$\psi_{it}=A_{3t}P_{t}^{-1}(p_{it}u_{it}+\overline{\mu}_{it}^{I})-(\overline{\mu}_{it}^{III}+\xi_{it}),$ (A.63)
where $A_{3t}=\mathbb{E}\overline{p}_{it}^{\prime}$ and $\overline{\mu}_{it}^{III}=\mathbb{E}\left[\left.b_{1t,V}\left(V_{jt},W_{j}\right)q_{jt}^{\prime}Q_{t}^{-1}q_{it}v_{ji,t}\right|\mathcal{I}_{i,t}\right]$.
Let
$\Omega_{3}:=\mathbb{\mathbb{V}}\left(\dfrac{1}{T}\sum_{t=1}^{T}\psi_{it}\right)=\dfrac{1}{T^{2}}\sum_{s=1}^{T}\sum_{t=1}^{T}\text{Cov}(\psi_{it},\psi_{is}).$ (A.64)
Following the same argument as for $\widehat{b}(x)$, I have
$\sqrt{n}\,\Omega_{3}^{-1/2}(\widehat{\overline{b}}-\overline{b})=\sqrt{n}\,\Omega_{3}^{-1/2}\left[\dfrac{1}{T}\sum_{t=1}^{T}(\widehat{b}_{t}^{\text{APE}}-b_{t}^{\text{APE}})\right]=\dfrac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\Omega_{3}^{-1/2}\,\dfrac{1}{T}\sum_{t=1}^{T}\psi_{it}\right)+o_{p}(1)$ (A.65)
and the asymptotic normality follows by the Lindeberg-Feller CLT:
$\sqrt{n}\,\Omega_{3}^{-1/2}(\widehat{\overline{b}}-\overline{b})\overset{d}{\longrightarrow}N(0,I).$ (A.66)
Finally, by (A.64), an estimator $\widehat{\Omega}_{3}$ for $\Omega_{3}$ is written as
$\widehat{\Omega}_{3}$ $=\dfrac{1}{T^{2}}\sum_{s=1}^{T}\sum_{t=1}^{T}\left(\dfrac{1}{n}\sum_{i=1}^{n}\widehat{\psi}_{it}\widehat{\psi}_{is}^{\prime}\right),$ (A.67)
where $\widehat{\psi}_{it}=\widehat{A}_{3t}\widehat{P}_{t}^{-1}(\widehat{p}_{it}\widehat{u}_{it}+\widehat{\overline{\mu}}_{it}^{I})-(\widehat{\overline{\mu}}_{it}^{III}+\widehat{\xi}_{it})$, $\widehat{A}_{3t}=n^{-1}\sum_{i}\widehat{\overline{p}}_{it}^{\prime}$, and $\widehat{\overline{\mu}}_{it}^{III}=n^{-1}\sum_{j}\widehat{b}_{1t,V}\left(\widehat{v}_{jt},w_{j}\right)q_{jt}^{\prime}\widehat{Q}^{-1}q_{it}\widehat{v}_{ji,t}.$ The consistency is established following the same argument as in deriving (A.62):
$\widehat{\Omega}_{3}/\Omega_{3}-1\overset{p}{\longrightarrow}0.$ (A.68)
Remark 5 (Degree of Basis Functions).
The three degrees $M_{1},M_{2},M_{3}$ correspond to the series dimensions used to estimate $V_{t}(x,z,w)$, $G_{t}(x,v,w)$, and $b_{t}(x)$, respectively. For consistency I require $M_{i}\to\infty$ while $M_{i}/n\to 0$, so that approximation bias and sampling variance both vanish. Under Assumption 5, the first-step estimator satisfies
$\text{MSE}(\widehat{V}_{t})=O\left(\underbrace{M_{1}^{1-2d_{1}/j_{1}}}_{\text{bias}}+\underbrace{M_{1}/n}_{\text{variance}}\right),$
where the extra factor $M_{1}$ in the bias term comes from the uniform-in-$x$ analysis with $\mathbf{1}\{X_{it}\leq x\}$ as the dependent variable (as in IN02). For steps 2–3 the analysis is analogous, with two additional points: (i) since the model implies $G_{t}$ is linear in $x$, I use the reduced basis $p^{M_{2}}(x,v,w)=x\otimes p^{m_{2}}(v,w)$, which reduces dimension ($M_{2}=m_{2}d_{X}$); and (ii) first-step errors propagate, so the error for $\widehat{b}_{t}(x)$ aggregates across steps
$\text{MSE}(\widehat{b}_{t}(x))=\left(M_{1}^{1-2d_{1}/j_{1}}+M_{1}/n\right)+\left(m_{2}^{-2d_{2}/j_{2}}+m_{2}/n\right)+\left(M_{3}^{-2d_{3}/j_{3}}+M_{3}/n\right).$
Sup-norm results additionally impose standard bounds on basis functions and derivatives (Assumption 6(b)), satisfied by common choices such as splines and power series. Finally, Assumption 7 imposes “not too fast” growth on $M_{i}$ for asymptotic normality; for instance, for $\widehat{\overline{b}}$ (which does not involve $M_{3}$), with splines it suffices that $M_{1}=o(n^{1/7})$ and $M_{2}=o(n^{1/7})$, which is feasible when $d_{1}/j_{1}\geq 4$ and $d_{2}/j_{2}\geq 4$.
A.2 Nonparametric Justification of the Index Exclusion Assumption
Next, I introduce a proposition that supplies sufficient conditions for Assumption 2 by adapting the exchangeability condition of Altonji and Matzkin (2005).
Proposition 1 (Sufficient Conditions for Assumption 2).
@@ A.2 Nonparametric Justification of the Index Exclusion Assumption
To begin with, I use the first condition $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ to establish the exchangeability condition
$f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}=f_{\left.A_{i}\right|\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}},$ (A.69)
where $\left(\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}\right)$ is any permutation of $\left({\bf X}_{i},{\bf Z}_{i}\right)$ in time. It is worth emphasizing that (A.69), which will be proved in the following, is different from Assumption 2.3 of Altonji and Matzkin (2005). In particular, Altonji and Matzkin (2005) assume an exchangeability condition involving both $A$ and $\eta_{t}$, i.e., $f_{\left.A_{i},\eta_{it}\right|{\bf X}_{i}}=f_{\left.A_{i},\eta_{it}\right|{\bf\widetilde{X}}_{i}},$ which effectively rules out time-varying endogeneity through the random coefficients because it requires the density of $\eta_{it}$ given $X_{it}=x_{it}$ to be the same as that given $X_{it}=x_{is}$ for any $s\neq t$. Instead, I prove (A.69) under a more primitive assumption of $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ that is compatible with time-varying endogeneity through the random coefficients.
Without loss of generality, I prove (A.69) for $T=2$ since any ordering of $\left\{1,\ldots,T\right\}$ for finite $T$ can be achieved via a finite number of pairwise permutations. Then, I prove that one can construct $W_{i}$ such that Assumption 2 holds. For simplicity of notation, I assume $X_{it}$ and $Z_{it}$ are both scalars and suppress $i$ subscripts in all variables. Thus, in this proof all subscripts of the variables denote the time period.
@@
$f_{A,\eta_{1},\eta_{2}}\left(a,h_{1},h_{2}\right)=f_{A,\eta_{1},\eta_{2}}\left(a,h_{2},h_{1}\right).$ (A.70)
@@
$=$ $\ f_{\left.X_{1},X_{2},A\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1},a\right|z_{2},z_{1}\right),$ (A.71)
where the first equality holds by the change of variables for $\eta_{1}$ and $\eta_{2}$, the second equality uses (A.70) and $Z\perp\left(A,\eta\right)$, the third equality holds by $X_{1}=g\left(z_{2},a,g^{-1}\left(x_{2},z_{2},a\right)\right)=x_{2}$ and $X_{2}=g\left(z_{1},a,g^{-1}\left(x_{1},z_{1},a\right)\right)=x_{1}$, and the last equality uses the fact that the product of derivatives of inverse functions equals one.
Given (A.71), I integrate AA out in (A.71) to obtain
$f_{\left.X_{1},X_{2}\right|Z_{1},Z_{2}}\left(\left.x_{1},x_{2}\right|z_{1},z_{2}\right)$ $=f_{\left.X_{1},X_{2}\right|Z_{1},Z_{2}}\left(\left.x_{2},x_{1}\right|z_{2},z_{1}\right),$ (A.72)
@@
$=$ $\ f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{2},x_{1},z_{2},z_{1}\right),$ (A.73)
where the second equality uses (A.71) and (A.72).
The rest of the proof follows the same argument as in Section 2.2 of Altonji and Matzkin (2005). Specifically, I show that for any on-support $a$, the conditional density $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by a function of the form $f_{\left.A\right|W}\left(\left.a\right|w\right)$. To construct $W$, I follow Weyl (1939) and let $\varphi_{1}\left(u\right)=\sum_{t=1}^{T}u_{t}$, $\varphi_{2}\left(u,v\right)=\sum_{t\neq s}^{T}u_{t}v_{s}$, …, $\varphi_{T}\left(u,v,\ldots,w\right)=\sum_{t\neq s\neq...\neq k}^{T}u_{t}v_{s}\ldots w_{k}$, where $u,v,\ldots,w$ are generic $T\times 1$ vectors. Then, I substitute each column vector of ${\bf D}_{i}=\left({\bf X}_{i},{\bf Z}_{i}\right)$ for each of the arguments $u,v,\ldots,w$ (repetitions included) to construct $W_{i}$. See Chapter II.3 of Weyl (1939) for details of the construction and its proof.
By condition (b) of Proposition 1, the supports of $X$ and $Z$ are compact. By condition (c) of Proposition 1, $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}$ is continuous in $\left(X_{1},X_{2},Z_{1},Z_{2}\right)$. Therefore, by the Stone-Weierstrass theorem, there exists a function $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}$ that is a polynomial in $\left(X_{1},X_{2},Z_{1},Z_{2}\right)$ over a compact set with the property that for any fixed $\delta$ that is arbitrarily close to 0,
$\max_{x_{1},x_{2}\in\mathcal{X},z_{1},z_{2}\in\mathcal{Z}}\left|f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)-f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)\right|\leq\delta.$ (A.74)
@@
denote the simple average of $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ over all $T!$ (here $T=2$) unique permutations of $\left(x_{t},z_{t}\right)$, and similarly for $\overline{f}_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}^{w}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$. By (A.73),